# Exploring relaxation dynamics in warm dense plasmas by tailoring non-thermal electron distributions with a free electron laser

### v3

@@ III.2 Electron-electron and electron-ion collisions
In the preceding section, we showed how we can track the relaxation of the electron distribution function by varying the pulse duration of the XFEL. This process, however, includes the contributions of both e-i and e-e collisions, both of which thermalize the distribution, but in different ways. To better understand their relative roles we shall now examine how the electron distribution function changes in time due to these two types of collisions. The simulation corresponds to 10 $\mu$m thin Al foil irradiated by an x-ray pulse with a photon energy of 2000 eV, pulse duration of 0.5 fs, and a peak intensity of $10^{19}$ Wcm-2. The first derivative of the electron distribution function $\frac{\partial f}{\partial t}$ is shown at peak intensity in Fig. 6, split into the two contributions. We note that the e-e contribution to $\frac{\partial f}{\partial t}$ is negative at the centres of the non-thermal peaks, but positive in the outer wings. These collisions tend to transfer relatively small amounts of energy, and their effect is thus gradually to broaden the non-thermal peaks. In contrast, the combined effect of inelastic collisions and the source term, which include both e-i collisions and photoionization and radiative recombination (PI/RR), manifest differently. Here, the contribution is positive at the centre of the peak, due to non-thermal electron generation via PI or Auger processes. In the surrounding areas, the contribution turns negative as collisional ionization and recombination redistribute the non-thermal electrons.
Figure 7: Contribution of e-i and e-e collisions to electron thermalization. The electron distribution function is shown in (a), and the corresponding emission spectrum in (b). Note how the e-e collisions primarily determine the shape of the emission spectrum, which is the main experimental observable.
The effect of collisional processes on the electron distribution and on the observed spectrum can be further explored by selectively deactivating them in the atomic kinetics simulations and in computing the emission spectrum. We show these results in Fig. 7. When only e-e collisions are active, we note that the shape of the non-thermal peaks in the distribution function remains similar to those calculated with the full simulation, with the only notable difference being their overall intensity. In turn, e-i collisions are critical to redistributing energy away from the non-thermal peaks across a wide energy range, as we observe in the region around 800 eV, far from the K-shell photoionization and Auger features. Here, e-e collisions play a negligible role in the overall equilibration. Similarly, e-i collisions only have a minor effect in determining the shape of the non-thermal peaks, which are, instead, predominantly broadened by e-e collisions.
@@
III.3 Effect of e-e relaxation time scales
We now explore how different electron relaxation timescales, particularly for e-e collisions, affect the spectral emission. These timescales are important in determining relaxation and transport in warm-dense plasmas, and such a comparison sheds light on how these processes are impacted by the accuracy of collisional models, often described in terms of a Coulomb Logarithm. We have already shown that the pulse duration of an XFEL largely determines the shape of the emission spectrum. Thus, we can imagine an experiment where multiple pulse durations are used to drive a system to the plasma state, and the emission spectrum is recorded. The electron relaxation timescale can then be determined by scanning across collisional models in the simulations until agreement with the experiment is reached.
The previous simulations assumed relaxation time determined by Huba’s version of the Coulomb Logarithm. As a first step, we could expand this by adding additional models as shown in Fig. 8, but we quickly see that at the solid densities we are interested in, many models suffer from unphysical divergences. In practice, simulations deploying CL models, such as CCFLY, tend to check for such divergences and replace the CL with a constant (often $\ln[\Lambda]=1$) once it starts approaching 0. So, rather than using different CL models to show the impact on relaxation timescales, we will scale the Huba relaxation time of $\sim 0.2$ fs instead.
@@ III.3 Effect of e-e relaxation time scales
Figure 8: Different Coulomb Logarithm models can lead to unphysically diverging relaxation times at high densities and low temperatures.
@@
Figure 9: Non-thermal features under different assumptions of e-e relaxation timescales. a). The relaxation is assumed to be 10 times as faster as that given by Huba’s formula. The non-thermal features are weak even for the shortest pulse duration. b). The relaxation time is 0.1 times that of Huba, and non-thermal features are prominent. Note that even the L shell RR has appeared for this group of simulations.
@@
We performed two sets of simulations where the elastic e-e collision operator in Eq. (2) is manually rescaled by a factor of 0.1 or 10. We again run simulations over a range of pulse durations and show the results for the emission spectra in Fig. 9. For all simulations, we have maintained the normal cross-section for all e-i collisions; these are not affected by any scaling factor. We are primarily interested in how the spectra change as a function of e-e collisional rates only. When the relaxation time is made 10 times faster than that of the Huba model, even the shortest simulated XFEL pulse of 0.1 fs cannot detect any noticeable non-thermal features in the emission spectrum apart from the Auger knee just above 3 keV. We note that at present, x-ray pulse durations shorter than this are not readily accessible at XFEL facilites [49]. The contrasting case where the relaxation time is made longer (2 fs) than the XFEL pulse duration is shown in Fig. 9b. Here, the nonthermal emission features are clearly visible. For the shortest 0.1 fs x-ray pulses, we observe a multitude of satellite peaks, including from L-shell recombination. As non-thermal electrons persist for longer, many of them can recombine at a later time when various configurations of ions with a K shell core hole have appeared, resulting in richer emission features.
@@ IV Conclusions
In this study, we have described how the relaxation time of free electrons in the warm dense matter regime can be inferred from the emission spectra created by tailored non-thermal electron distributions. By varying the pulse duration of an ultra-short-pulse XFEL beam, the thermalization process of non-thermal electrons can be observed and studied without the need for time-resolved measurements. For the non-thermal features to be prominent in the emission spectrum it is necessary for the x-ray pulse duration to be comparable in length to the typical e-e thermalization timescale. In warm-dense Al, this is expected to be on the order of 0.2 fs, but would be higher in mid- and high-Z materials driven at higher x-ray photon energies. The ability of modern XFELs to generate narrow-bandwidth, energetic attosecond pulses thus makes them ideally suited for such measurements. This work describes an approach that may soon be viable to explore directly the validity of the e-e collisional models, including those based on the Coulomb Logarithm framework, in describing equilibration and transport in warm dense plasmas.

### v4

@@ III.2 Electron-electron and electron-ion collisions
In the preceding section, we showed how we can track the relaxation of the electron distribution function by varying the pulse duration of the XFEL. This process, however, includes the contributions of both e-i and e-e collisions, both of which thermalize the distribution, but in different ways. To better understand their relative roles we shall now examine how the electron distribution function changes in time due to these two types of collisions. The simulation corresponds to 10 $\mu$m thin Al foil irradiated by an x-ray pulse with a photon energy of 2000 eV, pulse duration of 0.5 fs, and a peak intensity of $10^{19}$ Wcm-2. The first derivative of the electron distribution function $\frac{\partial f}{\partial t}$ is shown at peak intensity in Fig. 6, split into the two contributions. We note that the e-e contribution to $\frac{\partial f}{\partial t}$ is negative at the centres of the non-thermal peaks but positive in the outer wings. These collisions tend to transfer relatively small amounts of energy, and their effect is thus gradually to broaden the non-thermal peaks. In contrast, the combined effect of inelastic collisions and the source term, which include both e-i collisions and photoionization and radiative recombination (PI/RR), manifest differently. Here, the contribution is positive at the centre of the peak due to non-thermal electron generation via PI or Auger processes. In the surrounding areas, the contribution turns negative as collisional ionization and recombination redistribute the non-thermal electrons.
Figure 7: Contribution of e-i and e-e collisions to electron thermalization. The electron distribution function is shown in (a), and the corresponding emission spectrum in (b). The short-time relaxation of the electron spikes is driven by e-e collisions, while the overall equilibration across a larger energy range is primarily determined by e-i collisions.
The effect of collisional processes on the electron distribution and on the observed spectrum can be further explored by selectively deactivating them in the atomic kinetics simulations and in computing the emission spectrum. We show these results in Fig. 7. When only e-e collisions are active, we note that the shape of the non-thermal peaks in the distribution function remains similar to those calculated with the full simulation, with the only notable difference being their overall intensity. In turn, e-i collisions are critical to redistributing energy away from the non-thermal peaks across a wide energy range, as we observe in the region around 800 eV, far from the K-shell photoionization and Auger features. Here, e-e collisions play a negligible role in the overall thermalization. Similarly, e-i collisions only have a minor effect in determining the shape of the non-thermal peaks, which are, instead, predominantly broadened by e-e collisions.
@@
Figure 8: Different CL models can lead to unphysically diverging relaxation times at high densities (a) and low temperatures (b).
@@
Figure 9: Non-thermal features under different assumptions of e-e relaxation timescales. a). The relaxation is assumed to be 10 times as fast as that given by Huba’s formula. The non-thermal features are weak even for the shortest pulse duration. b). The relaxation rate is 0.1 times that of Huba, and non-thermal features are prominent. Note that even the L shell RR has appeared for this group of simulations. Note that similar to Fig. 5, offsets are applied to different pulse durations to provide a better comparison
III.3 The effect of e-e relaxation timescales
We now explore how different electron relaxation timescales, particularly for e-e collisions, affect the spectral emission. These timescales are important in determining relaxation and transport in warm-dense plasmas, and such a comparison sheds light on how these processes are impacted by the accuracy of collisional models, often described in terms of a CL. We have already shown that the pulse duration of an XFEL largely determines the shape of the emission spectrum. Thus, we can imagine an experiment where multiple pulse durations are used to drive a system to the plasma state, and the emission spectrum is recorded. The electron relaxation timescale can then be determined by scanning across collisional models in the simulations until agreement with the experiment is reached.
The previous simulations assumed relaxation time determined by Huba’s version of CL. As a first step, we could expand this by adding additional models as shown in Fig. 8, but we quickly see that at the solid densities we are interested in, many models suffer from unphysical divergences. In practice, simulations deploying CL models, such as CCFLY, tend to check for such divergences and replace the CL with a constant (often $\ln[\Lambda]=1$) once it starts approaching 0. So, rather than using different CL models to show the impact on relaxation timescales, we will scale the Huba relaxation time of $\sim 0.2$ fs instead.
@@ III.3 The effect of e-e relaxation timescales
We performed two sets of simulations where the elastic e-e collision operator in Eq. (5) is manually rescaled by a factor of 0.1 or 10. We again run simulations over a range of pulse durations and show the results for the emission spectra in Fig. 9. For all simulations, we have maintained the normal cross-section for all e-i collisions; these are not affected by any scaling factor. We are primarily interested in how the spectra change as a function of e-e collisional rates only. When the relaxation time is made 10 times as fast as that of the Huba model, even the shortest simulated XFEL pulse of 0.1 fs cannot detect any noticeable non-thermal features in the emission spectrum apart from the Auger knee just above 3 keV. We note that at present, x-ray pulse durations shorter than this are not readily accessible at XFEL facilites [58]. The contrasting case where the relaxation time is made longer (2 fs) than the XFEL pulse duration is shown in Fig. 9b. Here, the nonthermal emission features are clearly visible. For the shortest 0.1 fs x-ray pulses, we observe a multitude of satellite peaks, including from L-shell recombination. As non-thermal electrons persist for longer, many of them can recombine at a later time when various configurations of ions with a K shell core hole have appeared, resulting in richer emission features. The distinction between panels a and b in Figure 9 shows how the different relaxation speed influences the spectral response to pulse durations scan and serve as a key indicator for measuring the relaxation time in warm dense matter. Moreover, if the relaxation time is close to that in Fig 9 b (much smaller than that predicted by the CL model with Huba’s formula), the difference between spectra is subtle, indicating the primary limitation of our proposed method.
@@ IV Conclusions
In this study, we have described how the relaxation time of free electrons in the warm dense matter regime can be inferred from the emission spectra created by tailored non-thermal electron distributions. By varying the pulse duration of an ultra-short-pulse XFEL beam, the thermalization process of non-thermal electrons can be observed and studied without the need for time-resolved measurements. For the non-thermal features to be prominent in the emission spectrum it is necessary for the x-ray pulse duration to be comparable in length to the typical e-e thermalization timescale. We have shown that the relaxation time is estimated as 0.2 fs using the Spitzer and Härm model. Because electrons with higher energies typically take longer to thermalize, we expect the thermalization time in mid or high-Z material to be longer than that in Al, though the relaxation time may be similar. The ability of modern XFELs to generate narrow-bandwidth, energetic attosecond pulses thus makes them ideally suited for such measurements. This work describes an approach that may soon be viable to explore directly the validity of the e-e collisional models, including those based on the CL framework, in describing equilibration and transport in warm dense plasmas.